# Strategy for Direct Detection of Chiral Phonons with Phase-Structured Free Electrons

### v1

@@ Abstract
where $\mathbf{R}_{0}$ allows for the displacement of the transition field interference pattern in the $xy$ plane relative to some arbitrarily placed origin. In close analogy to topological electromagnetic vorticies created by interfering free space [4, 5] and evanescent [6, 7, 8, 9, 10] electromagnetic waves with controlled geometric phase, the in-plane polarization of the transition field $\mathbf{E}^{0\parallel}_{Pk_{\parallel}\sigma\phi_{0}}(\mathbf{x})$ consists of a 2D periodic array of field vorticies. The presently discussed topological field textures are distinguished from those previously reported in that the transition field is sourced by the charge density associated with the probing free electron transition, leading to field texture periodicity on the $\sim$ Å length scale. For the case $P=3$ and $k_{\parallel}=\textrm{K,\,K}^{\prime}$, the phased momentum transfers have the same three-fold rotational reciprocal space symmetry as the K and $\textrm{K}^{\prime}$ phonons [12, 13], and the real space polarization array has hexagonal symmetry with unit cell dimensions matching those of the atomic crystal as presented in Fig. 2b for $\phi_{0}=0$. Superimposed above the 2D $|\mathbf{E}^{0\parallel}_{3k_{\parallel}\sigma\phi_{0}}(\mathbf{x}_{\parallel},z=0)|$ maps are polarization ellipses colored red (blue) to indicate anticlockwise (clockwise) polarization rotation. Reversing the winding direction of the plane wave phases by $\sigma=+1\leftrightarrow\sigma=-1$ leads to a positional shift of the field nodes relative to the atomic crystal sites. The transition fields shown in Fig. 2b,c exhibit spin-momentum locking in the sense that simultaneous exchange of $\sigma=+1\leftrightarrow\sigma=-1$ and $\phi_{0}=0\leftrightarrow\phi_{0}=\pi/3$, which is equivalent to $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$, yields the same spatial in-plane field intensity profile, with regions of clockwise and anticlockwise polarization rotation interchanged.
The ability of the identified free electron transitions to generate electric fields that mirror the spin texture and spin-momentum locking of the hBN chiral phonons makes it possible to leverage the former to directly probe the latter. Fig. 3a presents calculated transition rates $w_{\sigma}^{\phi_{0}}(\hbar\omega_{if},\mathbf{R}_{0})$ as a function of $\hbar\omega_{if}$ and $\phi_{0}$ from Eq. (1) with $\mathbf{E}^{0}_{fi,\mathbf{k}+\mathbf{G}\kappa}$ from Eq. (4) for transitions $|P=3,k_{\parallel}=\textrm{K},\sigma=\pm 1,\phi_{0}\rangle\to|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$ with $\mathbf{R}_{0}=-a_{\textrm{BN}}/2$. Red and blue traces correspond to $\sigma=+1$ and $\sigma=-1$, respectively. At this $\mathbf{R}_{0}$ and $\phi_{0}=0$, as depicted in Fig. 2b, the transition field either drives both sublattices equally with circular polarizations of opposite handedness ($\sigma=+1$), or the B sublattice is situated at a node of the driving field while the N sublattice is at an antinode of the circularly polarized driving field ($\sigma=-1$). Consequently, as can be rationalized from the real-space phonon mode motions depicted in Fig. 1c, only the modes at 119 meV and 141 meV appear in Fig. 3a for $\phi_{0}=0$, where each of the momentum recoils is associated with one of the three BZ K points. As $\phi_{0}$ varies, the fixed-magnitude momentum recoils leave the BZ boundary and the features in the spectra shown in Fig. 3a evolve according to the hBN eigenmode dispersion along the circle of fixed $k_{\parallel}$. At $\phi_{0}=\pi/3$, the momentum recoils are associated with the three BZ $\textrm{K}^{\prime}$ points, as shown in Fig. 2c, and the spectra in Fig. 3a are identical to those at $\phi_{0}=0$ with $\sigma=+1\leftrightarrow\sigma=-1$ as anticipated from behavior of the phonon modes shown in Fig. 1c under $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$.
@@
Figure 3: Chiral phonon discrimination with phase shaped electrons. a) Calculated transition rates for incoming pinwheel states transitioning to the on-axis plane wave state as a function of loss energy and $\phi_{0}$ for an hBN monolayer with $\mathbf{R}_{0}=-a_{\textrm{BN}}/2$. Red (blue) traces indicate $\sigma=+1$ ($\sigma=-1$). b) Evolution of $\textrm{CD}_{\sigma}^{\phi_{0}}(\hbar\omega_{\lambda})$ as a function of $\phi_{0}\in[0,\pi]$ for selected loss energies and hBN and graphene targets. For hBN, $\mathbf{R}_{0}=-a_{\textrm{BN}}/2$ for $\hbar\omega_{\lambda}=111$ meV and $\mathbf{R}_{0}=+a_{\textrm{BN}}/2$ for $\hbar\omega_{\lambda}=135$ meV. $\mathbf{R}_{0}=-a_{\textrm{CC}}/2$ for the graphene mode at $\hbar\omega_{\lambda}=144$ meV. c) Spatial maps of $\textrm{CD}_{\sigma}^{\phi_{0}}(\hbar\omega_{if},\mathbf{R}_{0})$ for 119 meV (hBN), 135 meV (hBN), and 144 meV (graphene) with $\phi_{0}=0$ (upper) and $\phi_{0}=\pi/3$ (lower), and $\mathbf{R}_{0}$ varied throughout the unit cell. All calculations include an empirical damping rate of $\eta=1$ meV/$\textrm{meV}/\hbar$ at zero temperature and a vacuum background.
This signature of chiral phonon spin-momentum locking can be quantified by defining
$\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})=2\frac{w^{\phi_{0}}_{+}(\hbar\omega_{if},\mathbf{R}_{0})-w^{\phi_{0}}_{-}(\hbar\omega_{if},\mathbf{R}_{0})}{\textrm{max}\big{[}w^{\phi_{0}}_{+}(\hbar\omega_{if},\mathbf{R}_{0})\big{]}+\textrm{max}\big{[}w^{\phi_{0}}_{-}(\hbar\omega_{if},\mathbf{R}_{0})\big{]}},$ (5)
where $\textrm{max}\big{[}w^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})\big{]}$ denotes the maximum loss rate over all $\phi_{0}$ values. For fixed $\mathbf{R}_{0}$ and $\phi_{0}$, $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ quantifies the differential inelastic scattering rate associated with ±$\pm 1$ units of PAM exchanged between the probing free electron and target at a given loss energy. The blue trace in Fig. 3b shows $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0}=-a_{\textrm{BN}}/2)$ for the hBN monolayer with $\phi_{0}\in[0,\pi]$ and loss energy 119 meV. Mirroring the sublattice spin associated with the chiral phonon [12], $\textrm{CD}^{\phi_{0}}_{\sigma}\sim 0$, except where $\textrm{CD}^{\phi_{0}}_{\sigma}=\pm 1$ at $\phi_{0}=m({\pi}/{3})$ $(m\in\mathbb{Z})$ due to complete suppression of loss rates for either $\sigma=\pm 1$ at these $\phi_{0}$ values. This behavior, evident also in Fig. 3a, is a consequence of PAM conservation during the inelastic scattering process [12, 27]. Since $|P=3,k_{\parallel},\sigma=\pm 1,\phi_{0}\rangle\to|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$ is associated with $\Delta\textrm{PAM}=-\sigma$, these transitions couple only to phonons with $\textrm{PAM}=+\sigma$. The spin-momentum locking of the hBN phonon modes exhibited in Fig. 1c under $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$, therefore, underlies the sign change of $\textrm{CD}^{\phi_{0}=m(\pi/3)}_{\sigma}$ in Fig. 3b as a function of $m$.
Similarly, when $\mathbf{R}_{0}=+a_{\textrm{BN}}/2$, the transition field profiles in Figs. 2b,c are shifted vertically upwards relative to the fixed atomic sites such that the N sublattice sites reside at nodes of the transition field, while the B sites reside at antinodes of $\mathbf{E}^{0\parallel}_{3k_{\parallel}\sigma\phi_{0}}(\mathbf{x})$ with circular polarization states for $\sigma=+1$ and $\phi_{0}=0$ [64]. This transition field couples to the K/$\textrm{K}/\textrm{K}^{\prime}$ phonon modes at $135$ meV (pink trace) associated with precession of the B sublattice sites (Fig. 1c). The opposite sign of the PAM associated with the 119 meV and 135 meV modes at both K and $\textrm{K}^{\prime}$ result in the opposite signs of $\textrm{CD}^{\phi_{0}=m(\pi/3)}_{\sigma}$ for these two loss energies at each $m$ as shown in Fig. 3b. Unlike in hBN where the lack of spatial inversion symmetry leads to the non-degenerate K/$\textrm{K}/\textrm{K}^{\prime}$ phonons with $\textrm{PAM}=\pm 1\hbar$ at 119 meV and 135 meV, spatial inversion symmetry in graphene leads to the degeneracy of the phonon modes associated with circular motion on either sublattice at K/$\textrm{K}/\textrm{K}^{\prime}$, and a net $\textrm{PAM}=0\hbar$ at $\hbar\omega_{if}=144$ meV (see SM [64]). Therefore, as a point of comparison, Fig. 3b includes $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if}=144{\textrm{ meV}},\mathbf{R}_{0}=-a_{\textrm{CC}}/2)$ for a graphene monolayer (green trace). As anticipated based upon the PAM conservation argument invoked above, $\textrm{CD}^{\phi_{0}=m(\pi/3)}_{\sigma}=0$ for graphene at 144 meV as depicted in Fig. 3b. See SM [64] for additional results pertaining to the evolution of $\textrm{CD}^{\phi_{0}}_{\sigma}$ for hBN and graphene.
In contrast to the optical techniques previously employed to measure chiral phonons, atomic scale spatial information regarding sublattice spin can be accessed using the proposed inelastic free electron scattering strategy. Varying $\mathbf{R}_{0}$, e.g., by spatially translating the (monolayer) sample relative to the fixed widefield periodic array of transition field vorticies, allows the various local in-plane polarization states of the $\mathbf{E}_{3k_{\parallel}\sigma\phi_{0}}^{0\parallel}(\mathbf{x})$ site-space unit cell to be scanned across the atomic positions, modulating the electron-chiral-phonon coupling. We emphasize that the proposed detection scheme does not require careful a priori alignment of the transition field relative to the atomic crystal sites. Instead, it is required that the sign of $\sigma=\pm 1$ can be reversed keeping $\mathbf{R}_{0}$ fixed such that $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ is well defined. Spatial maps of $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ are presented in Fig. 3c for the hBN and graphene K/$\textrm{K}/\textrm{K}^{\prime}$ modes considered in Fig. 3b, while other in-plane modes are presented in the SM [64]. In the case of hBN, the spatial maps clearly reflect the spatial inversion symmetry breaking and the localization of the sublattice spin (phonon motion) onto one or the other sublattices as exhibited in Fig. 1c. The spatial maps of $\textrm{CD}^{\phi_{0}}_{\sigma}(\hbar\omega_{if}=144\textrm{ meV},\mathbf{R}_{0})$ for graphene are odd with respect to $y$ reflection due to spatial inversion symmetry. The ability of this inelastic electron scattering strategy to encode Å-scale spatial information directly related to the rotational motion of the sublattice sites constitutes a significant advantage over previous chiral phonon characterization techniques.
@@ Acknowledgements.
All work was supported by the U.S. Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences (BES), Materials Sciences and Engineering Division under Award No. DOE BES DE-SC0022921.

### v2

@@ Abstract
Here, $\mathbf{R}_{0}$ dictates the displacement of the transition field interference pattern in the $xy$ plane relative to some arbitrarily placed origin, which we take as the center of a real-space unit cell. In analogy to topological electromagnetic vorticies created by interfering free space Gao et al. (2020); Shen et al. (2021) and evanescent Tsesses et al. (2018); Du et al. (2019); Dai et al. (2020); Davis et al. (2020); Ghosh et al. (2023) electromagnetic waves with controlled geometric phase, the in-plane polarization of the transition field $\mathbf{E}^{0\parallel}_{Pk_{\parallel}\sigma\phi_{0}}(\mathbf{x})$ consists of a 2D periodic array of field vorticies. The presently discussed topological field textures are distinguished from those previously reported in that the transition field is sourced by the charge density associated with the probing free electron transition, leading to field texture periodicity on the $\sim$ Å length scale. For the case $P=3$ and $k_{\parallel}=\textrm{K,\,K}^{\prime}$, the phased momentum transfers have the same three-fold rotational reciprocal space symmetry as the K and $\textrm{K}^{\prime}$ phonons Zhang and Niu (2015); Zhu et al. (2018), and the real space polarization array has hexagonal symmetry with unit cell dimensions matching those of the atomic crystal as presented in Fig. 2b for $\phi_{0}=0$. Superimposed above the 2D $|\mathbf{E}^{0\parallel}_{3k_{\parallel}\sigma\phi_{0}}(\mathbf{x}_{\parallel},z=0)|$ maps are polarization ellipses colored red (blue) to indicate anticlockwise (clockwise) polarization rotation. In close analogy to the situation encountered for valley phonons in 2D atomic crystals with hexagonal symmetry, the in-plane transition field has well-defined PAM $l_{fi}^{i}\in\{-1,0,+1\}$ relative to any of the three distinct $C_{3}^{i}$ $(i=1,2,3)$ rotation centers marked with colored dots in Figs. 2b,c SM_ . Reversing the winding direction of the plane wave phases by $\sigma=+1\leftrightarrow\sigma=-1$ leads to a spatial shift of the field nodes relative to the atomic crystal sites. Consequently, $l_{fi}^{i}$ does not depend on the value of $\sigma$, and $l_{fi}^{i}$ is tabulated in Fig. 2d for $\phi_{0}=0$. The transition fields shown in Fig. 2b,c exhibit spin-momentum locking in the sense that simultaneous exchange of $\sigma=+1\leftrightarrow\sigma=-1$ and $\phi_{0}=0\leftrightarrow\phi_{0}=\pi/3$, which is equivalent to $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$, yields the same spatial in-plane field intensity profile, with regions of clockwise and anticlockwise polarization rotation interchanged. This property reflects the fact that $l^{i}_{fi}\to-l^{i}_{fi}$ when $\phi_{0}=0\to\pi/3$ SM_ .
@@
Figure 3: Chiral phonon discrimination with phase shaped electrons. a) Configuration tables specifying the identities of co-located atomic crystal and transition field $C_{3}$ axes for $\mathbf{R}_{0}=\big{\{}\big{(}0,\pm a/2\big{)},\big{(}\sqrt{3}a/2,0\big{)}\big{\}}$, $\sigma=\pm 1$, and $\phi_{0}=\{0,\pi/3\}$. Circles at the vertices of each triangle represent B, N, and center (cent) atomic crystal positions, while circle color indicates which of the transition field $C_{3}^{i}$ axes resides at each site. The integer to the lower right of each diagram indicates $l^{\text{cent}}_{fi}$. b) Calculated transition rates $w^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ for incoming pinwheel states transitioning to the on-axis plane wave state as a function of loss energy and $30\leq\phi_{0}\leq\pi/$ 3 for an hBN monolayer with $\mathbf{R}_{0}=\big{(}0,\pm a/2\big{)}$. Red (blue) traces indicate $\sigma=+1$ ($\sigma=-1$). Integers adjacent to spectral features at K/$\text{K}/\text{K}^{\prime}$ indicate the $l^{\text{cent}}_{fi}$ from (a) relevant to each peak. c) Same as (b), but for monolayer graphene. All calculations assume a vacuum background environment and zero temperature. An empirical phonon damping rate of $\eta=1$ meV/$\textrm{meV}/\hbar$ is included.
The ability of the identified free electron transitions to generate electric fields that mirror the spin texture and spin-momentum locking of the hBN chiral phonons makes it possible to leverage the former to directly probe the latter. In particular, spatially scanning the transition field relative to the atomic crystal by changing $\mathbf{R}_{0}$ leads to either alignment, or varying degrees of misalignment, of the three $C_{3}$ axes of the transition field with those of the atomic crystal. When the two groups of $C_{3}$ axes are aligned, the PAM of both the transition field and the phonons can be simultaneously well-defined relative to a common origin, and coupling occurs between transition and phonon displacement fields with equal PAM as required by PAM conservation. Conversely, when the transition field and atomic crystal $C_{3}$ axes are misaligned, the PAMs of the phonon branches cease to be well-defined relative to the $C_{3}$ axis of the probe, and exchange of intrinsic and extrinsic PAM occurs.
When $\mathbf{R}_{0}\in\{(0,\pm a/2),(\sqrt{3}a/2,0)\}$, the transition field $C_{3}^{i}$ and atomic crystal rotation center positions coincide, with the particular value of $i$ at each site dictated by the other parameters characterizing the transition field, i.e., $\sigma=\pm 1$ and $\phi_{0}\in\{0,\pi/3\}$. The possible configurations of co-located $C_{3}$ axes are pictorially represented in Fig. 3a. For example, when $\mathbf{R}_{0}=(0,-a/2)$, $\sigma=+1$, and $\phi_{0}=0$, it is the $C_{3}^{1}$ transition field axis at the location of the $C_{3}^{\text{cent}}$ atomic crystal rotation center. Therefore, since $l^{1}_{fi}=0$, this transition field couples only to valley phonons with PAM $l^{\text{cent}}_{\lambda}=0$. However, for $\mathbf{R}_{0}=(0,-a/2)$, $\sigma=-1$, and $\phi_{0}=0$, $C_{3}^{2}$ coincides with $C_{3}^{\text{cent}}$, and phonons with PAM $l^{\text{cent}}_{\lambda}=+1$ are excited. This is precisely the behavior exhibited by the corresponding state- and energy-resolved transition rates $w^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ presented in Fig. 3b for hBN, which are evaluated using Eqs. (1) and (4). Red and blue traces correspond to $\sigma=+1$ and $\sigma=-1$, respectively. For $\mathbf{R}_{0}=(0,-a/2)$ and $\phi_{0}=0$, the red trace shows features at the energies of modes $\lambda=1$ and $\lambda=4$, which have PAM $l^{\text{cent}}_{\lambda}=0$, while the blue trace exhibits a single peak at mode $\lambda=2$ with $l^{\text{cent}}_{\lambda}=+1$. Meanwhile, for $\mathbf{R}_{0}=(0,+a/2)$ and $\phi_{0}=0$, in agreement with PAM conservation considerations encoded into the table in Fig. 3a, the $\sigma=+1$ and $\sigma=-1$ transition fields couple to valley phonons with PAM $l^{\text{cent}}_{\lambda}=-1$ and $l^{\text{cent}}_{\lambda}=0$, respectively. Using $l^{\text{cent}}_{fi}\to-l^{\text{cent}}_{fi}$ when $\phi_{0}=0\to\pi/3$, and $l^{\text{cent}}_{\lambda}\to-l^{\text{cent}}_{\lambda}$ when $\text{K}\to\text{K}^{\prime}$, the spectra in Fig. 3b at $\phi_{0}=\pi/3$ can likewise be rationalized in terms of PAM conservation. Therefore, given the ability to perform the identified pre- and post-selected pinwheel state measurements with control over $\mathbf{R}_{0}$, $\sigma$, and $\phi_{0}$, it possible to perform energy-, valley, and PAM-resolved inelastic free electron scattering measurements capable of accessing valley phonons at the BZ corners directly.
As a point of comparison, Fig. 3c considers monolayer graphene, where the presence of parity symmetry requires modes $\lambda=2,3$ to be degenerate at the K/$\text{K}/\text{K}^{\prime}$ valleys. Although eigenvectors $\bm{\xi}_{\lambda=2,3}(\mathbf{k}_{\parallel}={\bf\text{K}}/{\bf\text{K}}^{\prime})$ can be defined with well-defined parity, parity and PAM operators do not commute. Nevertheless, linear combinations of the parity states can be taken to construct eigenvectors in this degenerate space with well-defined PAM. These particular eigenvectors have the same symmetries as those encountered for hBN, i.e., $\bm{\xi}_{\lambda=2}$ and $\bm{\xi}_{\lambda=3}$ exhibit circular motion of carbon atoms on one or the other atomic sites, respectively, while the other atom remains stationary at its equilibrium position. As a consequence, at the K valleys, these degenerate modes have total PAM $l^{\text{cent}}_{\lambda=2}=+1$ and $l^{\text{cent}}_{\lambda=3}=-1$. Critically, Figs. 3b,c show that by comparing the PAM-resolved loss spectra measured at $\mathbf{R}_{0}=(0,\pm a/2)$ and $\sigma=\pm 1$, it is possible to directly assess non-degeneracy of valley phonon modes with PAM $l_{\lambda}^{\text{cent}}=\pm 1$ introduced in the absence of parity symmetry. Spatial maps of $w^{\phi_{0}}_{\sigma}(\hbar\omega_{if},\mathbf{R}_{0})$ for hBN and graphene and $\mathbf{R}_{0}$ varied throughout the real-space unit cell are presented in the SM SM_ .
@@ Acknowledgements.
We acknowledge Jordan A. Hachtel for valuable discussions and feedback on the strategy proposed. All work was supported by the U.S. Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences (BES), Materials Sciences and Engineering Division under Award No. DOE BES DE-SC0022921.